A rope is wrapped around the rim of a large uniform solid disk of mass 315 kg and radius 1.60 m. The horizontal disk is made to rotate by pulling on the rope with a constant force of 195 N. If the disk starts from rest what is its angular speed in rev/s at the end of 2.95 s?
Answer #1

SOLUTION :

Torque in this case

= 195(1.60) = 312 N m

The moment of inertia of a disk = 1/2(m)r^2

= 0.5(315)* (1.6)^2

= 403.2 kg m^2

Torque x time = moment of inertia x angular velocity

312 (2.95) = 403.2 W

W = 2.28 rad/s

or W = 2.28/2pi = 0.363 rev/s